hhanan:

You invest an initial $500 in an account that has an annual interest rate of 2%, compounded quarterly. How much money will you have in the account after 6 years? Round your answer to the nearest whole number.

hhanan:

i think it is 560

hhanan:

@jhonyy9 could you help me

Pa21rala:

It's gonna be tough but I'll try 🤔🤔🤔

mxddi3:

\[a=p_o(1+\frac{ r }{ n })^nt\] Po is the initial amount (in this case 500) r is the rate as a decimal (convert 2% to a decimal) n is the number of time compounded per year (quarterly=4) and t is the time in yrs (in this case it's 6 so our equation is: \[a=500(1+\frac{ .02 }{ 4 })^4(6)\] and then from there, you can put it into the calculator. are you rounding your answer to a certain decimal place?

mxddi3:

oh wait, whole number, ok i see, well can you do tht ^ and tell me what you get?

hhanan:

3646.51875

mxddi3:

woah there. ok so no. uhm, to start, we are going to do what is in the parentheses first. what is .02/4? add one to tht and tell me what you get.

mxddi3:

at the end, its saying to the power of 4(6) so the whole equation is to the 24th power.

ImBadKidTae:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mxddi3 at the end, its saying to the power of 4(6) so the whole equation is to the 24th power. \(\color{#0cbb34}{\text{End of Quote}}\) yep that's write

hhanan:

0.005

hhanan:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ImBadKidTae \(\color{#0cbb34}{\text{Originally Posted by}}\) @mxddi3 at the end, its saying to the power of 4(6) so the whole equation is to the 24th power. \(\color{#0cbb34}{\text{End of Quote}}\) yep that's write \(\color{#0cbb34}{\text{End of Quote}}\) you spelt it wrong

mxddi3:

ok yes so the new equation is: A=500(1+0.005)^24 this can be rewritten as: A=500(1.005)^24. now do tht in the calculator and tell me what you get.

hhanan:

563.579888

mxddi3:

yes, now what is 563.579888 rounded to the nearest whole number?

hhanan:

564

mxddi3:

yes, that's the answer. does it make sense? You were super close in your original answer.

hhanan:

thank you and it does make sense

mxddi3:

you're welcome <3

jhonyy9:

good job @mxddi3 congrats !